# Multi-Excitation Projective Simulation with a Many-Body Physics Inspired Inductive Bias

### v2

@@ Appendix C Further complexity results and proofs
Let $\{c_{m_{1}},\dots,c_{m_{x}}\}$ be any excitation configuration. Then we can use $h_{(1,1)}(c^{\prime},c_{m_{1}})$ and $h_{(1,1)}(c_{m_{1}},c^{\prime})$ for any clip $c^{\prime}$ to move back and forth between $\{c_{m_{1}},\dots,c_{m_{x}}\}$ and $\{c^{\prime},c_{m_{2}},\dots,c_{m_{x}}\}$ arbitrarily many times. ∎
In Proposition 14 we showed that the total number of trainable parameters is polynomial in the number of clips. As a consequence, also the number of transitions that we need to consider in each deliberation step scales polynomially. However, do we need to consider all trainable parameters for sampling a transition? The following proposition gives a tighter bound:
@@
Proof.
Compared to Proposition 14, this result can provide significant benefits if the number $x$ of excitations in the current configuration is small (note that $x\leq|V|$ always), but the number of clips $|V|$ is very large.
Next, we aim to prove Proposition 16, which provides an upper bound on the maximal deliberation time of Inductive Bias 2FF scaling exponentially with the depth $D$.
Proposition (Proposition 16).
Assume Inductive Bias 2FF for a many-body mePS agent with layers $(L_{1},\dots,L_{D})$. Then the deliberation time (i.e. the total number of random walk steps) is upper-bounded by $\prod_{j=1}^{D-1}(|L_{j}|+1)$.
@@
Since transitions within Segment 2 do not change the number of steps needed for Segment 1, we can empty $L_{D-1}$ in Segment 2 every time new excitations arrive, before continuing with emptying Segment 1. This ensures that every time Segment 1 moves excitations into Segment 2, the clips are empty and no excitations are discarded in layer $L_{D-1}$. If we did not always empty layer $L_{D-1}$ first, it would not affect the number of steps needed to empty Segment 1, but it may reduce the total number of transitions within Segment 2 because moving a new excitation into an already excited clip just discards the excitation. That is one less excitation to remove one-by-one. This argument shows that the worst random walk empties the deepest non-final layer first.
@@
The first product on the left-hand side is the upper bound on the transitions within Segment 1, and the second product expresses that for each transition within Segment 1, we also have up to $|L_{D-1}|$ transitions within Segment 2.
In Proposition 16, we did not consider low cutoffs for $IO$. This may raise the question of whether such cutoffs dodge the exponential complexity. However, in general this is not the case, as we will see in the following proposition:
@@
While there is a upper bound on the maximal deliberation time, it scales exponentially with the depth $D$. The expression $\prod_{k=j}^{D-1}o$ from the previous proposition shows that in general, an exponential (in $D$) maximal deliberation time is unavoidable. However, the proofs of these exponential bounds required that one first has to remove excitations from the deepest layers. This observation motivated us to introduce Inductive Bias 2SF.

### v3

@@ Appendix D Detailed Complexity Results and Proofs
Let $\{c_{m_{1}},\dots,c_{m_{x}}\}$ be any excitation configuration. Then we can use $h_{(1,1)}(c^{\prime},c_{m_{1}})$ and $h_{(1,1)}(c_{m_{1}},c^{\prime})$ for any atomic clip $c^{\prime}$ to move back and forth between $\{c_{m_{1}},\dots,c_{m_{x}}\}$ and $\{c^{\prime},c_{m_{2}},\dots,c_{m_{x}}\}$ arbitrarily many times. ∎
One important consequence of our feed-forward conditions is that they explicitly break such cycles. In that regard, we first establish that our Inductive Bias 2FF indeed leads to a finite upper bound on the deliberation time:
@@
Assume Inductive Bias 2FF for a layered many-body mePS agent with layers $(L_{1},\dots,L_{D})$. Then the deliberation time (i.e. the total number of random walk steps) is upper-bounded by $\prod_{j=1}^{D-1}(|L_{j}|+1)$.
@@
Since transitions within Segment 2 do not change the number of steps needed for Segment 1, we can empty $L_{D-1}$ in Segment 2 every time new excitations arrive, before continuing with emptying Segment 1. This ensures that every time Segment 1 moves excitations into Segment 2, the atomic clips are empty and no excitations are discarded in layer $L_{D-1}$. If we did not always empty layer $L_{D-1}$ first, it would not affect the number of steps needed to empty Segment 1, but it may reduce the total number of transitions within Segment 2 because moving a new excitation into an already excited atomic clip just discards the excitation. That is one less excitation to remove one-by-one. This argument shows that the worst random walk empties the deepest non-final layer first.
@@
The first product on the left-hand side is the upper bound on the transitions within Segment 1, and the second product expresses that for each transition within Segment 1, we also have up to $|L_{D-1}|$ transitions within Segment 2. ∎
While the upper bound in Proposition 18 is finite, it is exponential in the depth $D$. However, there exist scenarios in which there is also an exponential lower bound on the maximal deliberation time, which is the content of the following proposition.
@@
The expression $\prod_{k=j}^{D-1}o$ from the previous proposition shows that in general, an exponential (in $D$) maximal deliberation time is unavoidable. To get this exponential scaling, it is enough to consider examples with $IO=\{(1,2)\}$, despite the fact that this $IO$ only contains small $i$ and $o$. This shows that small values of $i$ and $o$ are not enough to prevent exponentially large random walk lengths.
However, one key point of the proofs is that for all these scenarios, we require $(i,o)$ with $o>i$ such that we can start an “avalanche” of excitations. This suggests that, if we choose our inductive bias on $IO$ such that the number of excitations cannot increase, we find a much better upper bound, which is essentially $\mathrm{width}\times\mathrm{depth}^{2}$:
Proposition 20.
Proof.
Furthermore, the proofs of the exponential lower bound on the worst-case deliberation times required that one first has to remove excitations from the deepest layers. This observation motivated us to introduce Inductive Bias 2SF, which enforces the opposite order, i.e. that shallow excitations need to be removed first. We now argue that this change leads to upper bounds in the maximal deliberation time that are linear in the width and depth of the layered ECM:
Proposition 21.
Proof.
Proposition 22.
Proof.
In Proposition 14 we showed that the total number of trainable parameters is polynomial in the number of atomic clips. As a consequence, also the number of transitions that we need to consider in each deliberation step scales polynomially. However, do we need to consider all trainable parameters for sampling a transition? The following proposition gives a tighter bound:
Proposition 23.
Proof.
Compared to Proposition 14, this result can provide significant benefits if the number $x$ of excitations in the current configuration is small (note that $x\leq|V|$ always), but the number of atomic clips $|V|$ is very large.